# ActiveInitSplat: How Active Image Selection Helps Gaussian Splatting

### v1

@@ Abstract
Gaussian splatting (GS) along with its extensions and variants provides outstanding performance in real-time scene rendering while meeting reduced storage demands and computational efficiency. While the selection of 2D images capturing the scene of interest is crucial for the proper initialization and training of GS, hence markedly affecting the rendering performance, prior works rely on passively and typically densely selected 2D images. In contrast, this paper proposes ‘ActiveInitSplat’, a novel framework for active selection of training images for proper initialization and training of GS. ActiveInitSplat relies on density and occupancy criteria of the resultant 3D scene representation from the selected 2D images, to ensure that the latter are captured from diverse viewpoints leading to better scene coverage and that the initialized Gaussian functions are well aligned with the actual 3D structure. Numerical tests on well-known simulated and real environments demonstrate the merits of ActiveInitSplat resulting in significant GS rendering performance improvement over passive GS baselines, in the widely adopted LPIPS, SSIM, and PSNR metrics.
@@ 1 Introduction
3D scene representation and novel view synthesis are fundamental tasks in computer vision with diverse applications including autonomous driving, robotics, and medical imaging just to name a few. The goal in these tasks is to develop models capable of fast scene rendering while maintaining high visual quality, minimizing storage requirements, and ensuring computational efficiency. Gaussian splatting (GS), first proposed in [12], along with its extended variants have sparked attention in the research community due to their efficient 3D scene representation via a set of anisotropic Gaussian functions, allowing for fast rendering with memory efficiency, and outstanding visual performance.
@@
Across different GS models, the selection of initial training images is known to be a crucial factor for the rendering performance, since they have a direct impact on the initialization and training of GS. However, existing GS frameworks hinge on a passively pre-selected and typically dense set of training images which ensures redundancy, along with increased processing and computational cost. Even in the few GS frameworks designed for efficient rendering with a limited number of images, passive selection may confine the expressiveness of the 3D scene representation, potentially leading to reduced scene coverage. This paper, in contrast, provides a general framework for active and strategic selection of a few yet informative training images captured at diverse viewpoints, to ensure effective GS rendering performance at a lower sampling cost.
@@
Figure 1: ActiveInitSplat in a nutshell. In contrast to existing Gaussian splatting (GS) methods that rely on passively (and possibly densely) collected 2D images of the scene of interest, ActiveInitSplat actively selects informative images to assist the initialization and training of GS. The active selection mechanism lies on optimizing a (black-box) function quantifying the quality of the resultant 3D point cloud from the selected images via density and occupancy criteria. The collected images are captured from diverse viewpoints, ensuring better scene coverage and facilitating the accurate alignment of the initialized Gaussian functions with the underlying 3D structure.
@@
While the idea of actively optimizing the training set of images seems natural and necessary, identifying a compact set of optimal camera views that lead to high-quality 3D scene representation, is fundamentally challenged by the need to guarantee diverse viewpoints without incurring redundancy. Specifically, (i) there is no explicit function in the existing literature that maps a certain set of images to the quality of the reconstructed 3D scene and, consequently, to GS rendering performance; (ii) even if such an explicit function were provided, it certainly would depend on the structure of the scene for which prior information is typically not available; and finally (iii) a sequential processing of multiple different sets of images to empirically assess GS rendering performance is impractical. In the face of these challenges, this paper proposes ‘ActiveInitSplat’, a novel active GS framework shown in Fig. 1, whose contributions can be summarized in the following aspects.
@@ 2.2 Gaussian splatting for novel view synthesis
Striving for high rendering quality and reduced training time compared to the ‘Mip-NeRF360’ NeRF-based approach, the recently developed Gaussian splatting (GS) methods provide principled frameworks to represent a 3D scene of interest using a set of Gaussian functions. To initialize these Gaussian functions so that they are aligned with the 3D structure, the initial GS approach in [12] relies on the SfM process to obtain a representation of the 3D scene. Yet, the SfM process may become inefficient when the number of training images is large, may introduce accumulated errors in the rendering process, and necessitates explicit camera pose knowledge [7]. To cope with the latter challenge, alternative pose-free GS approaches have been proposed in [8, 10] that rely on monocular depth estimation for 3D representation instead of SfM. However, they may still require extensive training time. Aiming at efficient training solutions, existing SfM-free GS methods [7, 23] capitalize on pre-trained models, including ‘DUSt3R’ [29] and its following ‘MASt3R’ [14], that provide geometry priors for densely sampled, pixel-aligned stereo points. Besides leveraging prior models, alternative SfM-free approaches have been proposed for NVS with efficiency; see e.g. [26, 31, 32, 30]. In all these approaches, the selection of the initial set of training images plays a significant role in the rendering performance. Despite their well-documented merits, all these approaches rely on a passively (and possibly densely) pre-selected set of training images, which may lead to redundant views along with less expressive scene representation and reduced scene coverage, hence motivating the need for active camera view selection to aid GS.
@@
Lying at the crossroads of robotics and computer vision, existing approaches including [15, 11] combine GS models with additional supervision from Voronoi graphs or voxel maps for active path planning, while others leverage GS-based scene representations for active mapping [5]. Here, we would like to emphasize that these prior works that integrate GS solutions in active mapping or active high-fidelity scene reconstruction for various robotic applications are distinct from the proposed ActiveInitSplat in the sense that the latter aims to find informative camera views for GS beforehand without need for incremental re-training and use of GS. In robotics settings, the recent work in [25] seeks instructive camera views for GS accounting for depth uncertainty-based criteria. Albeit interesting, it is tailored for robotic environments where the targeted scene is mainly focused on specific objects without showing generalizability to broader 3D scenes of interest. In addition, the proposed framework in [25] integrates ground-truth depth information from depth sensors, which may not be given in several practical settings. Identifying informative camera views for any 3D scene of interest beforehand without the need for re-training of GS during the selection process, and without need for any depth or prior scene information is still unexplored.
@@ 3 Preliminaries and Problem Formulation
where $\mathbf{c}(\mathbf{p})$ represents the color at any pixel $\mathbf{p}$ in the rendered image and $\mathbf{c}_{i}$ the color corresponding to the $i$th Gaussian that is either expressed via RGB values or spherical harmonics (SH) which are used to model view-dependent color variations. The function $G_{i}^{2D}(\cdot)$ denotes the projected 2D Gaussian function with mean and covariance given by
$\mathbf{m}_{i}^{2D}$ $=[u/b,l/b]^{\top},\;[u,l,z]^{\top}=\mathbf{L}\mathbf{W}[\mathbf{m}_{i},1]^{\top}$ (3a)
$\mathbf{\Sigma}_{i}^{2D}$ $=\mathbf{J}\mathbf{W}\mathbf{\Sigma}_{i}\mathbf{W}^{\top}\mathbf{J}^{\top}$ (3b)
where matrices $\mathbf{L},\mathbf{W}$ contain intrinsic and extrinsic camera parameters respectively, and $\mathbf{J}$ is the Jacobian matrix.
@@
All Gaussian parameters are trained so that the rendered images match the corresponding training ground-truth ones; i.e. $\{\mathbf{I}_{\tau}\}_{\tau=1}^{T}$. To ensure that the covariance matrix $\mathbf{\Sigma}_{i}$ of each Gaussian $i$ remains positive semi-definite during the optimization process, it is decomposed as $\mathbf{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$ where $\mathbf{R}_{i}$ is the $3\times 3$ diagonal scaling matrix and $\mathbf{S}_{i}$ represents the rotation matrix analytically expressed in quaternions. The selection of the initial set of camera poses $\{\mathbf{x}_{\tau}\}_{\tau=1}^{T}$ that provide the initial training images $\{\mathbf{I}_{\tau}\}_{\tau=1}^{T}$, is critical for the GS performance since they affect the initialization and training of GS. Nonetheless, the optimal selection of initial camera poses is a non-trivial task and thus most existing works rely on passive selection of training images. In the next section, we will outline our proposed method for active selection of informative camera poses $\{\mathbf{x}_{\tau}^{*}\}_{\tau=1}^{T}$ that can assist GS models.
@@
4 Active Camera View Selection for GS
Aiming at actively identifying the most informative camera poses $\{\mathbf{x}_{\tau}^{*}\}_{\tau=1}^{T}$ used for proper initialization and training of GS, the present work introduces a novel optimization framework that relies on density and occupancy criteria of the PC in the 3D space emanating from the selected camera views. The intuition is that if the images from the actively selected camera poses result in more dense PCs with better scene coverage compared to passive selection, then (i) the initialized Gaussians will better cover the scene, avoiding missing areas or gaps and (ii) the training images will be selected from diverse viewpoints, thus facilitating the optimization process of GS.
$O(\mathcal{P}_{C}(\mathbf{x};\mathbf{x}_{1},\ldots,\mathbf{x}_{t-1}))$ (5)
At slot $t+1$, the goal is to find $\mathbf{x}_{t+1}^{*}$ given the training set $\mathcal{D}_{t}:=\{(\mathbf{x}_{\tau},y_{\tau})\}_{\tau=1}^{t}$ for the GP surrogate model consisting of all previous $t$ camera poses collected at $\mathbf{X}_{t}:=\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{t}\right]^{\top}$ along with the corresponding (possibly noisy) black-box function evaluations $\mathbf{y}_{t}:=[y_{1},\ldots,y_{t}]^{\top}$. Here $y_{\tau}=r_{q}(\mathbf{x}_{\tau};\mathbf{x}_{1},\ldots,\mathbf{x}_{\tau-1})+n_{\tau}(\forall\tau)$ and $n_{\tau}\sim\mathcal{N}(0,\sigma_{n}^{2})$ denotes Gaussian noise uncorrelated across $\tau$. Surrogate modeling with GPs deems the sought objective function $r_{q}(\cdot)$ as random and assumes that it is drawn from a GP prior. This means that the random vector $\mathbf{r}_{t}:=[r_{q}(\mathbf{x}_{1}),r_{q}(\mathbf{x}_{2};\mathbf{x}_{1}),\ldots,r_{q}(\mathbf{x}_{t};\mathbf{x}_{1}\ldots\mathbf{x}_{t-1})]^{\top}$ containing all function evaluations up until slot $t$, is Gaussian distributed as $p(\mathbf{r}_{t}|\mathbf{X}_{t})=\mathcal{N}({\bf 0}_{t},{\bf K}_{t})(\forall t)$ where $\mathbf{K}_{t}$ is the covariance matrix whose $(\tau,\tau^{\prime})$ entry is $[{\bf K}_{t}]_{\tau,\tau^{\prime}}={\rm cov}(r_{q}(\mathbf{x}_{\tau};\mathbf{x}_{1},\ldots,\mathbf{x}_{\tau-1}),r_{q}(\mathbf{x}_{\tau}^{\prime};\mathbf{x}_{1},\ldots,\mathbf{x}_{\tau^{\prime}-1}))$ [21].
$\sigma_{t+1}^{2}({\bf x})$ $=\!\kappa(\mathbf{x},\mathbf{x},t+1,t+1)\!$
where it is intuitive that if the surrogate model captures well the objective function, then selecting $\mathbf{x}$ that maximizes the posterior mean in (7a) most likely provides the maximum of $r_{q}$ at this slot. Upon selecting $\mathbf{x}_{t+1}$, the corresponding (possibly noisy) function evaluation $y_{t+1}$ is obtained, and the training set for the surrogate model is augmented as $\mathcal{D}_{t+1}=\mathcal{D}_{t}\bigcup\{(\mathbf{x}_{t+1},y_{t+1})\}$ for the next iteration. This process is repeated up until slot $T$, as depicted in Fig. 2. Then, the informative set of the $T$ collected camera poses and views can be used for initialization and training of different GS models.

### v2

@@ Abstract
Gaussian splatting (GS) along with its extensions and variants provides outstanding performance in real-time scene rendering while meeting reduced storage demands and computational efficiency. While the selection of 2D images capturing the scene of interest is crucial for the proper initialization and training of GS, hence markedly affecting the rendering performance, prior works rely on passively and typically densely selected 2D images. In contrast, this paper proposes ‘ActiveInitSplat’, a novel framework for active selection of training images for proper initialization and training of GS. ActiveInitSplat relies on density and occupancy criteria of the resultant 3D scene representation from the selected 2D images, to ensure that the latter are captured from diverse viewpoints leading to better scene coverage and that the initialized Gaussian functions are well aligned with the actual 3D structure. Numerical tests on well-known simulated and real environments demonstrate the merits of ActiveInitSplat resulting in significant GS rendering performance improvement over passive GS baselines in both dense- and sparse-view settings, in the widely adopted LPIPS, SSIM, and PSNR metrics.
@@ 1 Introduction
3D scene representation and novel view synthesis are fundamental tasks in computer vision that have gained significant traction, driven by their applications in diverse domains such as autonomous driving, robotics, medical imaging, and surveillance, just to name a few. The goal in these tasks is to develop models capable of fast scene rendering while maintaining high visual quality, minimizing storage requirements, and ensuring computational efficiency. Gaussian splatting (GS), first proposed in [12], along with its extended variants have sparked attention in the research community due to their efficient 3D scene representation via a set of anisotropic Gaussian functions, allowing for fast rendering with memory efficiency, and outstanding visual performance.
@@
Across different GS models, the selection of initial training images is known to be a crucial factor for the rendering performance, since they have a direct impact on the initialization and training of GS; see e.g. [12]. However, existing GS frameworks hinge on a passively pre-selected and typically dense set of training images which ensures redundancy, along with increased processing and computational cost. Even in the few GS frameworks designed for efficient rendering with a limited number of images, passive selection may confine the expressiveness of the 3D scene representation, potentially leading to reduced scene coverage. This paper, in contrast, provides a general framework for active and strategic selection of a few yet informative training images captured at diverse viewpoints, to ensure effective GS rendering performance at a lower sampling cost.
@@
Figure 1: ActiveInitSplat in a nutshell. In contrast to existing Gaussian splatting (GS) methods that rely on passively (and possibly densely) collected 2D images of the scene of interest (upper part), ActiveInitSplat actively selects informative images to assist the initialization and training of GS (lower part). The active selection mechanism lies on optimizing a (black-box) function quantifying the quality of the resultant 3D point cloud from the selected images via density and occupancy criteria. The collected images are captured from diverse viewpoints, ensuring better scene coverage and facilitating the accurate alignment of the initialized Gaussian functions with the underlying 3D structure.
@@
Although the idea of actively optimizing the training set of images seems natural and necessary, identifying a compact set of optimal camera views that yields high-quality 3D scene representation is fundamentally challenged by the need to guarantee diverse viewpoints without incurring redundancy. Specifically, (i) there is no explicit function in the existing literature that maps a certain set of images to the quality of the reconstructed 3D scene and, consequently, to GS rendering performance; (ii) even if such an explicit function were provided, it certainly would depend on the structure of the scene for which prior information is typically not available; and finally (iii) a sequential processing of multiple different sets of images to empirically assess GS rendering performance is impractical. In the face of these challenges, this paper proposes ‘ActiveInitSplat’, a novel active GS framework shown in Fig. 1, whose contributions can be summarized in the following aspects.
@@ 2.2 Gaussian splatting for novel view synthesis
Striving for high rendering quality and reduced training time compared to the NeRF-based approaches, the recently developed Gaussian splatting (GS) methods provide principled frameworks to represent a 3D scene of interest using a set of Gaussian functions. To initialize these Gaussian functions so that they are aligned with the 3D structure, the initial GS approach in [12] relies on the SfM process to obtain a representation of the 3D scene. Yet, the SfM process may become inefficient when the number of training images is large, may introduce accumulated errors in the rendering process, and necessitates explicit camera pose knowledge [7]. To cope with the latter challenges, alternative pose-free GS approaches have been proposed in [8, 10] that rely on monocular depth estimation for 3D representation instead of SfM. However, they may still require extensive training time. Aiming at efficient training solutions, existing SfM-free GS methods [7, 26] capitalize on pre-trained models, including ‘DUSt3R’ [32] and its following ‘MASt3R’ [16], that provide geometry priors for densely sampled, pixel-aligned stereo points. Besides leveraging prior models, alternative SfM-free approaches have been proposed for NVS with efficiency; see e.g. [29, 37, 39, 36]. In all these approaches, the selection of the initial set of training images plays a significant role in the rendering performance. Despite their well-documented merits, all these approaches rely on a passively (and possibly densely) pre-selected set of training images, which may lead to redundant views along with less expressive scene representation and reduced scene coverage, hence motivating the need for active camera view selection to aid GS. Note that the discussion above summarizes key fundamental passive GS methods as representative examples. We omit the many other passive GS variants in the literature, as our main focus is on active GS approaches, which are outlined next.
@@
Lying at the crossroads of robotics and computer vision, existing approaches including [17, 11] combine GS models with additional supervision from Voronoi graphs or voxel maps for active path planning, while others leverage GS-based scene representations for active mapping [5]. Here, we would like to emphasize that these prior works that integrate GS solutions in active mapping or active high-fidelity scene reconstruction for various robotic applications are distinct from the proposed ActiveInitSplat in the sense that the latter aims to find informative camera views for GS beforehand without need for incremental re-training and use of GS. The recent works in [33, 35] have explored active view selection in GS through mutual information criteria between residuals and Gaussian parameters or via maximizing P-optimal information gain within a continuously trained GS model. However, these methods require frequent retraining of the GS map as new observations are incorporated or inherently assume an actively maintained GS model, necessitating iterative updates to maintain meaningful uncertainty estimates. In contrast, ActiveInitSplat introduces an active view selection framework that is agnostic to the underlying GS architecture and operates without the need for sequential GS model updates. In robotics settings, the recent work in [28] seeks instructive camera views for GS accounting for depth uncertainty-based criteria. Albeit interesting, it is tailored for robotic environments where the targeted scene is mainly focused on specific objects without showing generalizability to broader 3D scenes of interest. In addition, the proposed framework in [28] integrates ground-truth depth information from depth sensors, which may not be given in several practical settings. Identifying informative camera views for any 3D scene of interest beforehand without the need for re-training of GS during the selection process, and without need for any depth or prior scene information is still unexplored.
Note that while we acknowledge that part of the existing literature addresses active view selection for NeRF-based methods; see e.g., [23, 34, 13], GS models have recently demonstrated state-of-the-art performance in both 3D scene rendering quality and computational efficiency. Therefore, our work, and the related-work discussion, focus exclusively on active view selection for GS models. Notably, since the proposed active view selection strategy in ActiveInitSplat is independent of the subsequent stages of GS models, it can also be readily extended to NeRF-based frameworks.
@@ 3 Preliminaries and Problem Formulation
where $\mathbf{c}(\mathbf{p})$ represents the color at any pixel $\mathbf{p}$ in the rendered image and $\mathbf{c}_{i}$ the color corresponding to the $i$th Gaussian that is either expressed via RGB values or spherical harmonics (SH) which are used to model view-dependent color variations. The function $G_{i}^{2D}(\cdot)$ denotes the projected 2D Gaussian function; see further details e.g. in [12].
@@
All Gaussian parameters are trained so that the rendered images match the corresponding training ground-truth ones; i.e. $\{\mathbf{I}_{\tau}\}_{\tau=1}^{T}$. To ensure that the covariance matrix $\mathbf{\Sigma}_{i}$ of each Gaussian $i$ remains positive semi-definite during the optimization process, it is decomposed as $\mathbf{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$ where $\mathbf{S}_{i}$ is the $3\times 3$ diagonal scaling matrix and $\mathbf{R}_{i}$ represents the rotation matrix analytically expressed in quaternions. The selection of the initial set of camera poses $\{\mathbf{x}_{\tau}\}_{\tau=1}^{T}$ that provide the initial training images $\{\mathbf{I}_{\tau}\}_{\tau=1}^{T}$, is critical for the GS performance since they affect its initialization and training. Nonetheless, the optimal selection of initial camera poses is a non-trivial task and thus most existing works rely on passive selection of training images. In the next section, we will outline our proposed method for active selection of proper camera poses $\{\mathbf{x}_{\tau}^{*}\}_{\tau=1}^{T}$ capturing informative camera views $\{\mathbf{I}_{\tau}^{*}\}_{\tau=1}^{T}$ that can assist GS models.
4 Active Camera View Selection for GS
Aiming at actively identifying the most informative camera poses $\{\mathbf{x}_{\tau}^{*}\}_{\tau=1}^{T}$ used for proper initialization and training of GS, the present work introduces a novel optimization framework that relies on density and occupancy criteria of the PC in the 3D space emanating from the selected camera views. The intuition is that if the images from the actively selected camera poses result in more dense PCs with better scene coverage compared to passive selection, then (i) the initialized Gaussians will better cover the scene, avoiding missing areas or gaps and (ii) the training images will be selected from diverse viewpoints, thus facilitating the training process of GS.
$O(\mathcal{P}_{C}(\mathbf{x};\mathbf{x}_{1},\ldots,\mathbf{x}_{t-1}))$ (4)
Relying on the GP-based surrogate model, the goal at slot $t+1$ is to find $\mathbf{x}_{t+1}^{*}$ given the training set $\mathcal{D}_{t}:=\{(\mathbf{x}_{\tau},y_{\tau})\}_{\tau=1}^{t}$ consisting of all previous $t$ camera poses collected at matrix $\mathbf{X}_{t}:=\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{t}\right]^{\top}$ along with the corresponding (possibly noisy) black-box function evaluations/observations collected at the vector $\mathbf{y}_{t}:=[y_{1},\ldots,y_{t}]^{\top}$. Here $y_{\tau}=r_{q}(\mathbf{x}_{\tau};\mathbf{x}_{1},\ldots,\mathbf{x}_{\tau-1})+n_{\tau}(\forall\tau)$ and $n_{\tau}\sim\mathcal{N}(0,\sigma_{n}^{2})$ denotes Gaussian noise uncorrelated across $\tau$. Surrogate modeling with GPs deems the sought objective function $r_{q}(\cdot)$ as random and assumes that it is drawn from a GP prior. This means that the random vector $\mathbf{r}_{t}:=[r_{q}(\mathbf{x}_{1}),r_{q}(\mathbf{x}_{2};\mathbf{x}_{1}),\ldots,r_{q}(\mathbf{x}_{t};\mathbf{x}_{1}\ldots\mathbf{x}_{t-1})]^{\top}$ containing all function evaluations up until slot $t$, is Gaussian distributed as $p(\mathbf{r}_{t}|\mathbf{X}_{t})=\mathcal{N}({\bf 0}_{t},{\bf K}_{t})(\forall t)$ where $\mathbf{K}_{t}$ is the covariance matrix whose $(\tau,\tau^{\prime})$ entry is ${\bf K}_{t}]_{\tau,\tau^{\prime}}={\rm cov}(r_{q}(\mathbf{x}_{\tau};\mathbf{x}_{1},\ldots,\mathbf{x}_{\tau-1}),r_{q}(\mathbf{x}_{\tau}^{\prime};\mathbf{x}_{1},\ldots,\mathbf{x}_{\tau^{\prime}-1})) [$24].
$\sigma_{t+1}^{2}({\bf x})$ $=\!\kappa(\mathbf{x},\mathbf{x},t+1,t+1)\!$
where it is intuitive that if the surrogate model captures well the objective function, then selecting $\mathbf{x}$ that maximizes the posterior mean in (6a) most likely provides the maximum of $r_{q}$ at this slot. Upon selecting $\mathbf{x}_{t+1}$, the corresponding (possibly noisy) function evaluation $y_{t+1}$ is obtained, and the training set for the surrogate model is augmented as $\mathcal{D}_{t+1}=\mathcal{D}_{t}\bigcup\{(\mathbf{x}_{t+1},y_{t+1})\}$ for the next iteration. This process is repeated up until slot $T$, as shown in Fig. 2. Then, the informative set of the $T$ collected camera poses and views can be used for initialization and training of different GS models. The steps of ActiveInitSplat are outlined in Algorithm 1.
Remark 1. To the best of our knowledge, ActiveInitSplat is the first framework to introduce active view selection based on the 3D representation quality derived from the selected images, to facilitate GS initialization and training while remaining agnostic to subsequent GS steps. In this work, 3D representation quality is evaluated using density and occupancy criteria of the 3D PC; however, we would like to note that incorporating additional metrics could further enhance ActiveInitSplat’s performance. Potential extensions include connectivity metrics to detect gaps or discontinuities in otherwise continuous structures, and sharpness metrics to assess edge and detail clarity, among others. Integrating such criteria into the active selection process constitutes part of our future research agenda.
Algorithm 1 ActiveInitSplat algorithm.
1:Initialization: Images $\{\mathbf{I}_{0},\mathbf{I}_{1}\}$, camera poses $\{\mathbf{x}_{0},\mathbf{x}_{1}\}$
2:Set $y_{0}=0$
3:Compute $y_{1}=r_{q}(\mathbf{x}_{1};\mathbf{x}_{0})+n_{1}$ via (4).
4:Construct initial (training) set $\mathcal{D}_{1}=\{(\mathbf{x}_{i},y_{i})\}_{i=0}^{1}$
5:# Active camera view selection process
6:for $t=2,\ldots,T$ do
7:  Update parameters of kernel $\kappa$ maximizing the marginal likelihood.
8:  Update GP surrogate model $p(r_{q}({\bf x})|\mathcal{D}_{t})=\mathcal{N}(\mu_{t}({\bf x}),\sigma_{t}^{2}({\bf x}))$ via (6a) and (6b).
9:  Obtain next camera pose $\mathbf{x}_{t+1}$ via (7) and obtain corresponding image $\mathbf{I}_{t+1}$.
10:  Compute $y_{t+1}=r_{q}(\mathbf{x}_{t+1};\mathbf{x}_{0},\ldots,\mathbf{x}_{t})+n_{t+1}$ via (4).
11:  Augment training set $\mathcal{D}_{t+1}=\mathcal{D}_{t}\bigcup\{(\mathbf{x}_{t+1},y_{t+1})\}$.
12:end for
13:# Gaussian splatting initialization and training
14:Initialize Gaussian functions $\{G_{i}\}_{i=1}^{N}$ using $\mathcal{P}_{C}(\mathbf{x}_{T};\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1})$.
15:Train a certain GS model using $\{\mathbf{I}_{0},\ldots,\mathbf{I}_{T}\}$.
@@ 4.1 Gaussian process based surrogate model for black-box optimization
5 Numerical Evaluation
5.1 Implementation details
For demonstration purposes, the 3D representation $\mathcal{P}_{C}(\mathbf{x}_{t};\mathbf{x}_{1},\ldots,\mathbf{x}_{t-1})$ used for the black-box function evaluation $r_{q}(\mathbf{x}_{t};\mathbf{x}_{1}\ldots\mathbf{x}_{t-1})$ upon acquiring $\mathbf{x}_{t}$ at slot $t$, is estimated using COLMAP [25] similarly as in the traditional SfM-based GS model in [12]. However, we note that other alternative techniques for obtaining $\mathcal{P}_{C}(\cdot)$ from 2D images can be readily integrated into our ActiveInitSplat framework; see e.g. a related ablation study in 5.2.3. For the Gaussian process surrogate model, the kernel functions $\kappa_{s},\kappa_{t}$ used for computing $\kappa$ in (5) are selected to be RBF kernels, which is a typical choice across different GP-based tasks [24]. The kernel hyperparameters are estimated maximizing the marginal likelihood using the sklearn package and are re-fitted at the end of each iteration where a newly acquired $(\mathbf{x}_{t},y_{t})$ pair of camera-view and (noisy) function evaluation becomes available. Note that the kernel function has a significant role in the GP surrogate model and existing works aim to learn the proper kernel form; see e.g. [19, 31, 18, 6], that could be applied in our ActiveInitSplat but exceeds the scope of the current work.
The active selection process (lines 6–12 of Algorithm 1) was executed on an Intel Core i7-5930K CPU and a GeForce Tatan X GPU, while the GS training process (lines 14–15) was conducted on a Tesla V100-SXM2-16GB GPU hosted on Amazon servers. In experiments utilizing the SfM process for PC generation, the COLMAP software was employed, while for those leveraging the MASt3R pre-trained model in subsection 5.2.3, we used the implementation available at https://github.com/naver/mast3r. For the GS training and rendering process, we relied on the implementation from https://github.com/graphdeco-inria/gaussian-splatting.